Three friends all have ages that are consecutive integers.The sum of their ages is 180. What is the age of the oldest friend .
Answer by macston(5194)

Oldest friend=X+2; Middle friend=X+1; Youngest friend=X
X+(X+1)+(X+2)=180
3X+3=180 Subtract 3 from each side.
3X=177 Divide each side by 3
X=59 Youngest friend is 59.
X+2=59+2=61 ANSWER: Oldest friend is 61.
X+2=59+1=60 Middle friend is 60.
CHECK:
Sum is 180
59+60+61=180
180=180